Hi. I need help on how to solve the following problem. I've tried solving it many times but couldn't get the correct answer.
Find the partial arithmetic sum of the second set of 100 even terms. (102-200)

Answer by venugopalramana(3286)

Hi. I need help on how to solve the following problem. I've tried solving it many times but couldn't get the correct answer.
Find the partial arithmetic sum of the second set of 100 even terms. (102-200)
S=102+104+106+.................+300
N=100
A=102
D=2
TN=102+(100-1)2=102+198=300
SN=(N/2)[A+L]
S=(100/2)[102+300}=50*402=20100